# Invariants of persistence modules defined by order-embeddings

### v2

@@ 3. Invariants defined by order-embeddings
Let $\mathrm{E}=\mathrm{Emb}_{\mathcal{X}_{2}}^{\mathcal{P}}$ and consider the map $p:\mathrm{E}\times\mathrm{I}\to\{\leq_{\mathcal{P}}\}$ defined by
$p(f,U)=\left\{\begin{array}[]{ll}(f1,f2)&\textrm{if }U=\mathbb{I}_{\{1,2\}}\\ (f1,f1)&\textrm{if }U=\mathbb{I}_{\{1\}}\\ (f2,f2)&\textrm{if }U=\mathbb{I}_{\{2\}}\end{array}\right.$
Since the poset $\mathcal{P}$ is connected and has at least two elements, this map is surjective and induces an injective morphism $\mathbb{Z}^{\{\leq_{\mathcal{P}}\}}\to\mathbb{Z}^{\mathrm{E}\times\mathrm{I}}$. Let $\Phi:\mathbb{Z}^{\mathrm{E}\times\mathrm{I}}\to\mathbb{Z}^{\mathrm{E}\times\mathrm{I}}$ defined by
$\Phi_{f}:=\begin{pmatrix}0&1&-1\\ 0&0&1\\ 1&0&-1\end{pmatrix}\textrm{ for any }f\in\mathrm{E}.$
@@
Since $\mathcal{X}_{2}$ has three indecomposables, the set $\mathrm{E}=\mathrm{Emb}_{\mathcal{X}_{2}}^{\mathcal{P}}$ has three times the cardinality of $\{<_{\mathcal{P}}\}$, where $\{<_{\mathcal{P}}\}:=\{(a,b)\in\mathcal{P}^{2}|a<b\}$. This is in general strictly greater than |$|\{\leq_{\mathcal{P}}\}|=|\{<_{\mathcal{P}}\}|+|\mathcal{P}|$, so the invariant $\mathbf{mult}_{\mathcal{X}_{2},\mathcal{P}}$ is not surjective. We will see in Section 5 a way to reduce this redundant information.
@@
Let $\mathcal{P}$ be a connected poset such that every element lies in a chain of length $\geq 3$. Let $\mathcal{X}_{3}$ be the poset $\{1\to 2\to 3\}$. Then the invariants $\operatorname{\mathbf{rk}}_{\mathcal{P}}$ and $\mathbf{mult}_{\mathcal{X}_{3},\mathcal{P}}$ are equivalent.
@@
There are $6$ indecomposable objects in $\operatorname{\mathrm{mod}}\mathcal{X}_{3}$ which are
@@
Then one computes
$\left(\mathbf{mult}_{\mathcal{X}_{3},\mathcal{P}}(M)\right)_{(f,U)}=\left\{\begin{array}[]{ll}\dim M_{f1}-\operatorname{\mathrm{rk}}M_{f1\to f2}&\textrm{if }U=\mathbb{I}_{\{1\}}\\ \dim M_{f2}-\operatorname{\mathrm{rk}}M_{f1\to f2}-\operatorname{\mathrm{rk}}M_{f2\to f3}+\operatorname{\mathrm{rk}}M_{f1\to f3}&\textrm{ if }U=\mathbb{I}_{\{2\}}\\ \dim M_{f3}-\operatorname{\mathrm{rk}}M_{f2\to f3}&\textrm{ if }U=\mathbb{I}_{\{3\}}\\ \operatorname{\mathrm{rk}}M_{f1\to f2}-\operatorname{\mathrm{rk}}M_{f1\to f3}&\textrm{ if }U=\mathbb{I}_{\{1,2\}}\\ \operatorname{\mathrm{rk}}M_{f2\to f3}-\operatorname{\mathrm{rk}}M_{f1\to f3}&\textrm{ if }U=\mathbb{I}_{\{2,3\}}\\ \operatorname{\mathrm{rk}}M_{f1\to f3}&\textrm{ if }U=\mathbb{I}_{\{1,2,3\}}\end{array}\right.$
Now for $a<b$ in $\mathcal{P}$, by hypothesis there exists $c\in\mathcal{P}$ with $c<a$ or with $b<c$. Therefore, there exists $f\in\mathrm{Emb}_{\mathcal{X}_{3}}^{\mathcal{P}}$ such that $f(1)=a$ and $f(2)=b$, or such that $f(2)=a$ and $f(3)=b$. Using this embedding $f$, one can easily check the equivalence. ∎
@@
$\mathcal{P}=$$a$$b$$c$$d$
one has $|\mathrm{Emb}_{\mathcal{X}_{3}}^{\mathcal{P}}|=2$, so $|\mathrm{Emb}_{\mathcal{X}_{3}}^{\mathcal{P}}\times\mathrm{I}|=12$. However the rank of the invariant is $9=|\{\leq_{\mathcal{P}}\}|$. One can check the following $3$ relations
$\mathbf{mult}(\mathbb{I}_{\{b,c,d\}})-\mathbf{mult}(\mathbb{I}_{\{b,c\}})-\mathbf{mult}(\mathbb{I}_{\{b,d\}})+\mathbf{mult}(\mathbb{I}_{\{b\}})=0;$
@@
where $M$ denotes the indecomposable $\mathcal{P}$-module with dimension vector $(1,2,1,1)$ and $\mathbf{mult}=\mathbf{mult}_{\mathcal{X}_{3},\mathcal{P}}$.
@@
Now for $f\in\mathrm{Emb}_{\mathcal{X}^{\prime}_{3}}^{\mathcal{P}}$, denote by $\operatorname{\mathrm{rk}}M_{f1\to f2\oplus f3}$ the rank of the map $M_{f1}\rightarrow M_{f2}\oplus M_{f3}$ induced by $M_{f1\to f2}$ and $M_{f1\to f3}$.
@@
$\left(\mathbf{mult}_{\mathcal{X}^{\prime}_{3},\mathcal{P}}(M)\right)_{(f,U)}=\left\{\begin{array}[]{ll}\dim M_{f1}-\operatorname{\mathrm{rk}}M_{f1\to f2}-\operatorname{\mathrm{rk}}M_{f1\to f3}+\operatorname{\mathrm{rk}}M_{f1\to f2\oplus f3}&\textrm{ if }U=\mathbb{I}_{\{1\}}\\ \dim M_{f2}-\operatorname{\mathrm{rk}}M_{f1\to f2}&\textrm{ if }U=\mathbb{I}_{\{2\}}\\ \dim M_{f3}-\operatorname{\mathrm{rk}}M_{f1\to f3}&\textrm{ if }U=\mathbb{I}_{\{3\}}\\ \operatorname{\mathrm{rk}}M_{f1,f2}-\operatorname{\mathrm{rk}}M_{f1\to f2\oplus f3}&\textrm{ if }U=\mathbb{I}_{\{1,2\}}\\ \operatorname{\mathrm{rk}}M_{f1\to f3}-\operatorname{\mathrm{rk}}M_{f1\to f2\oplus f3}&\textrm{ if }U=\mathbb{I}_{\{1,3\}}\\ \operatorname{\mathrm{rk}}M_{f1\to f2\oplus f3}&\textrm{ if }U=\mathbb{I}_{\{1,2,3\}}\end{array}\right.$
Let $a<b$ be in $\mathcal{P}$. Then by hypothesis, there exists an embedding $f:\mathcal{X}^{\prime}_{3}\to\mathcal{P}$ with $f(1)=a$ and $f(2)=b$. Then we have
$\left(\operatorname{\mathbf{rk}}_{\mathcal{P}}(M)\right)_{(a,b)}=\left(\mathbf{mult}_{\mathcal{X}^{\prime}_{3},\mathcal{P}}(M)\right)_{\left(f,\mathbb{I}_{\{1,2\}}\right)}-\left(\mathbf{mult}_{\mathcal{X}^{\prime}_{3},\mathcal{P}}(M)\right)_{\left(f,\mathbb{I}_{\{1,2,3\}}\right)}.$
For $a\in\mathcal{P}$, by convexity, there exists $f\in\mathrm{Emb}_{\mathcal{X}^{\prime}_{3}}^{\mathcal{P}}$ with either $f(1)=a$ or with $f(2)=a$.
@@
    Note that here again, the cardinality of $\mathrm{Emb}_{\mathcal{X}^{\prime}_{3}}^{\mathcal{P}}\times\mathrm{I}$ is strictly greater than the rank of the invariant $\mathbf{mult}:=\mathbf{mult}_{\mathcal{X}^{\prime}_{3},\mathcal{P}}$. For instance if $\mathcal{P}$ is the poset
    $\mathcal{P}=$$a$$b$$c$$d$
    then $|\mathrm{Emb}_{\mathcal{X}^{\prime}_{3}}^{\mathcal{P}}|=2$ so $|\mathrm{Emb}_{\mathcal{X}^{\prime}_{3}}^{\mathcal{P}}\times\mathrm{I}|=12$, but one can check that the rank of $\mathbf{mult}$ is $10$. We have the relations
@@
    where $M$ is the indecomposable $\mathcal{P}$ module with dimension vector $(1,2,1,1)$.
@@
    We note that this weaker property is not sufficient in order to deduce the fact that $\mathbf{mult}_{\mathcal{X}_{3}^{\prime},\mathcal{P}}$ is strictly finer than $\operatorname{\mathbf{rk}}_{\mathcal{P}}$. Indeed consider the case where $\mathcal{P}$ is given as above. This poset satisfies the weaker property above, but there is no embedding $f:\mathcal{X}^{\prime}_{3}\to\mathcal{P}$ with $f(1)=a$ and $f(2)=b$. As mentioned above, the family $\left\{\mathbf{mult}(\mathbb{I}_{\{a\}}),\mathbf{mult}(\mathbb{I}_{\{b\}}),\mathbf{mult}(\mathbb{I}_{\{a,b\}})\right\}$ is not free, but one easily checks that the family $\left\{\operatorname{\mathbf{rk}}_{\mathcal{P}}(\mathbb{I}_{\{a\}}),\operatorname{\mathbf{rk}}_{\mathcal{P}}(\mathbb{I}_{\{b\}}),\operatorname{\mathbf{rk}}_{\mathcal{P}}(\mathbb{I}_{\{a,b\}})\right\}$ is free. Thus the invariant
    $\mathbf{mult}_{\mathcal{X}_{3}^{\prime},\mathcal{P}}$ is not finer than $\operatorname{\mathbf{rk}}_{\mathcal{P}}$, and in fact the two invariants are incomparable in this case. Similar examples occur when $\mathcal{X}_{3}^{\prime}$ is replaced with $\mathcal{X}_{3}^{\prime\prime}$.
The condition on $\mathcal{P}$ in Proposition 3.12 may seem strong. For instance, it is not satisfied when $\mathcal{P}$ is an $n\times m$ grid, $\mathcal{P}=\{1,2,\ldots,n\}\times\{1,2,\ldots,m\}$, since the ordered pairs between vertices of the form $(i,m)$ are not in the image of an order-embedding of $\mathcal{X}_{3}^{\prime}$. However, one can enlarge the poset $\mathcal{P}$ into one where the condition is satisfied, and then restrict to modules with support in $\mathcal{P}$:
@@
Let $\mathcal{P}=\{1,2,\ldots,n\}\times\{1,2,\ldots,m\}$ be an $n\times m$ grid with $n,m>1.$ Then the invariant $\mathbf{mult}_{\mathcal{X}^{\prime}_{3},\mathcal{P}}$ (resp. $\mathbf{mult}_{\mathcal{X}^{\prime\prime}_{3},\mathcal{P}}$) is strictly finer than the rank invariant $\operatorname{\mathbf{rk}}_{\mathcal{P}}$.
@@
$\mathcal{P}^{\prime}=$$(1,1)$$(2,1)$$(3,1)$$(4,1)$$(5,1)$$(1,2)$$(2,2)$$(3,2)$$(4,2)$$(1,3)$$(2,3)$$(3,3)$$(1,4)$$(2,4)$$(1,5)$
Clearly every pair $a<b\in\mathcal{P}^{\prime}$ is in the image of an order-embedding $\mathcal{X}_{3}\to\mathcal{P}^{\prime}$, and therefore by Proposition 3.12, the invariant $\mathbf{mult}_{\mathcal{X}^{\prime}_{3},\mathcal{P}^{\prime}}$ is strictly finer than the rank invariant $\operatorname{\mathbf{rk}}_{\mathcal{P}^{\prime}}$. The fact that the invariants $\mathbf{mult}_{\mathcal{X}^{\prime}_{3},\mathcal{P}^{\prime}}$ and $\operatorname{\mathbf{rk}}_{\mathcal{P}^{\prime}}$ are not equivalent is shown in the proof of Proposition 3.12 by embedding one single copy of $\mathcal{X}_{3}\to\mathcal{P}^{\prime}$, and this can be realized within the boundaries of $\mathcal{P}$ as soon as $n,m>1.$

### v3

@@ 3. Invariants defined by order-embeddings
Let $\mathrm{E}=\mathrm{Emb}_{\mathcal{X}_{2}}^{\mathcal{P}}$ and consider the map $p\mathrel{\mathop{\ordinarycolon}}\mathrm{E}\times\mathrm{I}\to\{\leq_{\mathcal{P}}\}$ defined by
$p(f,U)=\left\{\begin{array}[]{ll}(f2,f2)&\textrm{if }U=\mathbb{I}_{\{2\}}\\ (f1,f2)&\textrm{if }U=\mathbb{I}_{\{1,2\}}\\ (f1,f1)&\textrm{if }U=\mathbb{I}_{\{1\}}\end{array}\right$.
Since the poset $\mathcal{P}$ is connected and has at least two elements, this map is surjective and induces an injective morphism $p^{*}\mathrel{\mathop{\ordinarycolon}}\;\mathbb{Z}^{\{\leq_{\mathcal{P}}\}}\to\mathbb{Z}^{\mathrm{E}\times\mathrm{I}}$. Let $\Phi\mathrel{\mathop{\ordinarycolon}}\mathbb{Z}^{\mathrm{E}\times\mathrm{I}}\to\mathbb{Z}^{\mathrm{E}\times\mathrm{I}}$ defined by the matrix $\Phi_{f}$ above for any $f\in\mathrm{E}$ in the basis of $\mathbb{Z}^{\mathrm{E}\times\mathrm{I}}$ induced by the basis $(\mathbb{I}_{\{2\}},\mathbb{I}_{\{1,2\}},\mathbb{I}_{\{1\}})$ of $\mathbb{Z}^{I}$.
@@
Since $\mathcal{X}_{2}$ has three indecomposables, the set $\mathrm{E}\times I=\mathrm{Emb}_{\mathcal{X}_{2}}^{\mathcal{P}}\times\operatorname{\mathrm{Ind}}\mathcal{X}_{2}$ has three times the cardinality of $\{<_{\mathcal{P}}\}$, where $\{<_{\mathcal{P}}\}\mathrel{\mathop{\ordinarycolon}}=\{(a,b)\in\mathcal{P}^{2}|a<b\}$. This is in general strictly greater than $|\{\leq_{\mathcal{P}}\}|=|\{<_{\mathcal{P}}\}|+|\mathcal{P}|$, so the invariant $\mathbf{mult}_{\mathcal{X}_{2},\mathcal{P}}$ is not surjective. We will see in Section 5 a way to reduce this redundant information.
@@
Let $\mathcal{P}$ be a connected poset such that any two comparable elements lie in a chain of length $3$. Let $\mathcal{X}_{3}$ be the poset $\{1\to 2\to 3\}$. Then the invariants $\operatorname{\mathbf{rk}}_{\mathcal{P}}$ and $\mathbf{mult}_{\mathcal{X}_{3},\mathcal{P}}$ are equivalent.
@@
There are $6$ indecomposable objects in $\operatorname{\mathrm{mod}}\mathcal{X}_{3}$ which we label by
@@
One computes easily that
[…]
For comparable elements $a<b$ in $\mathcal{P}$, by hypothesis there exists $c\in\mathcal{P}$ with $c<a$, with $b<c$ or with $a<c<b$. Therefore, there exists $f\in\mathrm{Emb}_{\mathcal{X}_{3}}^{\mathcal{P}}$ with $a$ and $b$ in its image. Using this embedding $f$, one can easily check the equivalence. ∎
@@
$\mathcal{P}=$aabbccdd
[…]
$\mathbf{mult}(\mathbb{I}_{\{b,c,d\}})_{(f,U)}=\left\{\begin{array}[]{ll}1&\textrm{ for }U=\mathbb{I}_{\{2,3\}}\\ 0&\textrm{else}\end{array}\right$.
$\mathbf{mult}(\mathbb{I}_{\{b,c\}})_{(f,U)}=\left\{\begin{array}[]{ll}1&\textrm{ for $(f,U)=(f_{c},\mathbb{I}_{\{2,3\}})$ or $(f_{d},\mathbb{I}_{\{2\}})$}\\ 0&\textrm{else}\end{array}\right$.
and
$\mathbf{mult}(\mathbb{I}_{\{b\}})_{(f,U)}=\left\{\begin{array}[]{ll}1&\textrm{ for $U=\mathbb{I}_{\{2\}}$}\\ 0&\textrm{else}\end{array}\right$.
Hence we observe the following relation
$\mathbf{mult}(\mathbb{I}_{\{b,c,d\}})-\mathbf{mult}(\mathbb{I}_{\{b,c\}})-\mathbf{mult}(\mathbb{I}_{\{b,d\}})+\mathbf{mult}(\mathbb{I}_{\{b\}})=0.$
Similar computations give
@@
where $M$ denotes the indecomposable $\mathcal{P}$-module with dimension vector $(1,2,1,1)$. One further may check that the family in $K_{0}^{\rm sp}(\mathcal{P})$
$([\mathbb{I}_{\{b,c,d\}}\oplus\mathbb{I}_{\{b\}}]-[\mathbb{I}_{\{b,c\}}\oplus\mathbb{I}_{\{c,d\}}],\ [\mathbb{I}_{\{a,b,c,d\}}\oplus\mathbb{I}_{\{a,b\}}]-[\mathbb{I}_{\{a,b,c\}}\oplus\mathbb{I}_{\{a,b,d\}}],\ [M\oplus\mathbb{I}_{\{b\}}]-[\mathbb{I}_{\{a,b\}}\oplus\mathbb{I}_{\{b,c\}}\oplus\mathbb{I}_{\{b,d\}}])$
forms a basis of the kernel of $\mathbf{mult}\mathrel{\mathop{\ordinarycolon}}K_{0}^{\rm sp}(\mathcal{P})\to\mathbb{Z}^{\mathrm{E}\times\mathrm{I}}$.
@@
Now for $f\in\mathrm{Emb}_{\mathcal{X}^{\prime}_{3}}^{\mathcal{P}}$, denote by $\operatorname{\mathrm{rk}}M_{f1\to f2\oplus f3}$ the rank of the map
[…]
@@
[…]
Let $a<b$ be in $\mathcal{P}$. Then by hypothesis, there exists an embedding $f\mathrel{\mathop{\ordinarycolon}}\mathcal{X}^{\prime}_{3}\to\mathcal{P}$ with $f(1)=a$ and $f(2)=b$. This yields
$\left(\operatorname{\mathbf{rk}}_{\mathcal{P}}(M)\right)_{(a,b)}=\left(\mathbf{mult}_{\mathcal{X}^{\prime}_{3},\mathcal{P}}(M)\right)_{\left(f,\mathbb{I}_{\{1,2\}}\right)}+\left(\mathbf{mult}_{\mathcal{X}^{\prime}_{3},\mathcal{P}}(M)\right)_{\left(f,\mathbb{I}_{\{1,2,3\}}\right)}.$
For $a\in\mathcal{P}$, by connectedness, there exists $f\in\mathrm{Emb}_{\mathcal{X}^{\prime}_{3}}^{\mathcal{P}}$ with either $f(1)=a$ or with $f(2)=a$.
@@
    Note that here again, the cardinality of $\mathrm{Emb}_{\mathcal{X}^{\prime}_{3}}^{\mathcal{P}}\times\mathrm{I}$ is strictly greater than the rank of the invariant $\mathbf{mult}\mathrel{\mathop{\ordinarycolon}}=\mathbf{mult}_{\mathcal{X}^{\prime}_{3},\mathcal{P}}$. For instance if $\mathcal{P}$ is the poset
    $\mathcal{P}=$aabbccdd
    then $|\mathrm{Emb}_{\mathcal{X}^{\prime}_{3}}^{\mathcal{P}}|=2$ so $|\mathrm{Emb}_{\mathcal{X}^{\prime}_{3}}^{\mathcal{P}}\times\mathrm{I}|=12$, but one can check that the rank of $\mathbf{mult}$ is $10$. Denoting $f_{a}$ (resp. $f_{b}$) the embedding $\mathcal{X}\to\mathcal{P}$ with $f1=a$ (resp. $f1=b$), then one computes
    $\mathbf{mult}(M)_{(f,U)}=\left\{\begin{array}[]{ll}1&\textrm{if $(f,U)=(f_{a},\mathbb{I}_{\{1,2,3\}})$, $(f_{b},\mathbb{I}_{\{1,2\}})$ or $(f_{b},\mathbb{I}_{\{1,3\}})$}\\ 0&\textrm{else}\end{array}\right$.
    where $M$ is the indecomposable $\mathcal{P}$-module with dimension vector $(1,2,1,1)$.
    We therefore have the relations
@@
    that define a basis of the kernel of $\mathbf{mult}\mathrel{\mathop{\ordinarycolon}}K_{0}^{\rm sp}(\mathcal{P})\to\mathbb{Z}^{\mathrm{E}\times\mathrm{I}}$ as in Remark 3.11.
@@
    We note that this weaker property is not sufficient in order to deduce the fact that $\mathbf{mult}_{\mathcal{X}_{3}^{\prime},\mathcal{P}}$ is strictly finer than $\operatorname{\mathbf{rk}}_{\mathcal{P}}$. Indeed consider the case where $\mathcal{P}$ is given as above. This poset satisfies the weaker property above, but there is no embedding $f\mathrel{\mathop{\ordinarycolon}}\mathcal{X}^{\prime}_{3}\to\mathcal{P}$ with $f(1)=a$ and $f(2)=b$. As mentioned above, the family $\left\{\mathbf{mult}(\mathbb{I}_{\{a\}}),\mathbf{mult}(\mathbb{I}_{\{b\}}),\mathbf{mult}(\mathbb{I}_{\{a,b\}})\right\}$ is not free, but one easily checks that the family $\left\{\operatorname{\mathbf{rk}}_{\mathcal{P}}(\mathbb{I}_{\{a\}}),\operatorname{\mathbf{rk}}_{\mathcal{P}}(\mathbb{I}_{\{b\}}),\operatorname{\mathbf{rk}}_{\mathcal{P}}(\mathbb{I}_{\{a,b\}})\right\}$ is free. Thus the invariant $\mathbf{mult}_{\mathcal{X}_{3}^{\prime},\mathcal{P}}$ is not finer than $\operatorname{\mathbf{rk}}_{\mathcal{P}}$, and in fact the two invariants are incomparable in this case. Similar examples occur when $\mathcal{X}_{3}^{\prime}$ is replaced with $\mathcal{X}_{3}^{\prime\prime}$.
The condition on $\mathcal{P}$ in Proposition 3.12 may seem strong. For instance, it is not satisfied when $\mathcal{P}$ is an $n\times m$ grid, $\mathcal{P}=\{1,2,\ldots,n\}\times\{1,2,\ldots,m\}$, since the ordered pairs between vertices of the form $(i,m)$ are not in the image of an order-embedding of $\mathcal{X}_{3}^{\prime}$.
@@
Let $\mathcal{P}=\{1,2,\ldots,n\}\times\{1,2,\ldots,m\}$ be an $n\times m$ grid with $n,m>1.$ Then there exists an embedding $\mathcal{P}\to\mathcal{P}^{\prime}$ and an invariant $\overline{\mathbf{mult}}_{\mathcal{X}^{\prime}_{3},\mathcal{P}^{\prime},\mathcal{P}}$ (resp. $\overline{\mathbf{mult}}_{\mathcal{X}^{\prime\prime}_{3},\mathcal{P}^{\prime},\mathcal{P}}$) which is strictly finer than the rank invariant $\operatorname{\mathbf{rk}}_{\mathcal{P}}$.
@@
$\mathcal{P}^{\prime}=$(1,1)(1,1)(2,1)(2,1)(3,1)(3,1)(4,1)(4,1)(5,1)(5,1)(1,2)(1,2)(2,2)(2,2)(3,2)(3,2)(4,2)(4,2)(1,3)(1,3)(2,3)(2,3)(3,3)(3,3)(1,4)(1,4)(2,4)(2,4)$(1,5)$
Clearly every pair $a<b\in\mathcal{P}^{\prime}$ is in the image of an order-embedding $\mathcal{X}^{\prime}_{3}\to\mathcal{P}^{\prime}$, and therefore by Proposition 3.12, the invariant $\mathbf{mult}_{\mathcal{X}^{\prime}_{3},\mathcal{P}^{\prime}}$ is strictly finer than the rank invariant $\operatorname{\mathbf{rk}}_{\mathcal{P}^{\prime}}$. The fact that the invariants $\mathbf{mult}_{\mathcal{X}^{\prime}_{3},\mathcal{P}^{\prime}}$ and $\operatorname{\mathbf{rk}}_{\mathcal{P}^{\prime}}$ are not equivalent is shown in the proof of Proposition 3.12 by embedding one single copy of $\mathcal{X}^{\prime}_{3}\to\mathcal{P}^{\prime}$, and this can be realized within the boundaries of $\mathcal{P}$ as soon as $n,m>1.$